# It’s Not Easy Being Green: On the Energy Efficiency of Programming Languages

### v1

@@ 1. Introduction
The acceleration of climate change due to use of fossil fuels has driven an increased focus on efforts to decrease both the energy consumption and carbon footprint of computer systems (Hanafy et al., 2024b, a; Radovanovic et al., 2021). In 2018, an estimated 1% of total global energy consumption was attributed to datacenters alone (Masanet et al., 2020). Modern machine learning workloads—especially training—can generate hundreds of tons of CO2 emissions (Strubell et al., 2019). For example, Meta reports that training the Llama2 large language model generated an estimated 539 tons of CO2 emissions (GenAI, 2023).
@@
These studies have received wide attention, both in academic and industrial circles. They have been collectively cited over 500 times per Google Scholar (as of September 2024). The results—especially the ranking of programming languages—have had an unusually visible impact in industry, and are routinely quoted on social networks and in blog posts. As Figure 1 illustrates, the rankings have been cited by executives and engineers from Amazon (Miller and Lerche, 2022, 2021; Miller, 2022; Vogels, 2023), Intel (Mulhall and Collins, 2022), SAP (Muthaiah, 2022), and other companies (Suárez et al., 2021) to argue for business decisions and to advocate for a shift in programming languages with an eye towards sustainability. Often, these rankings have been harnessed to support the adoption of Rust, which ranks as one of the most energy-efficient languages while providing safety guarantees that languages like C and C++ lack.
@@
Our model specifically identifies the number of active cores and memory activity as the key contributors to power draw and consequently energy consumption. It highlights the importance of both programming language and application implementations: language implementation decisions can add runtime cost in the form of garbage collection or just-in-time warmup, while application implementations dictate the level of parallelism and overall performance of the program. These varying language and application implementation characteristics are the primary drivers of the anomalies we identify in prior work.
Our experimental approach incorporates several methodological improvements, correcting technical errors in prior work that led to negative or otherwise incorrect energy readings. We present an improved methodology based on hardware performance counter data that makes it possible to accurately quantify average core usage and account for memory activity. After controlling for confounds like varying CPU utilization by forcing benchmarks to run on a single core, and by normalizing by time, we conclude that energy consumed is directly proportional to runtime, and independent of the choice of programming language. These results suggest that, to minimize energy consumption, programmers should focus primarily on optimizing performance.
@@ 2.1. Overview
The line of work that is the point of departure for this paper explores the relationship between the choice of programming language and energy efficiency (Couto et al., 2017; Pereira et al., 2017, 2021). These papers rank programming languages based on energy consumption, runtime, and memory usage, and attempt to find associations between these metrics. In the remainder of this paper, we refer to these papers collectively as “Pereira et al.”.
Table 1. Partial results from Pereira et al. (§2.1), showing average runtime and energy consumption for the languages discussed in this paper, normalized to C. Boldface denotes anomalous results that Sections 2.2, 4.2, and 4.5 address.
@@
In essence, these studies compare the runtime, energy consumption and memory usage of benchmark implementations in 27 various programming languages. Table 1 presents some of the main results from Pereira et al., limited to the languages discussed in this paper, with anomalous results highlighted in boldface (Section 2.2 discusses this selection in detail). Their experiments leverage Intel’s Running Average Power Limit (RAPL) interface to measure energy consumption, and GNU time or Python’s memory_profiler to measure peak or “total” memory usage, respectively.
The programs used for comparison are from the Computer Language Benchmark Game (CLBG) (Gouy, 2024), a corpus of small benchmark implementations in various languages (22 to 287 lines of code). The most recent paper in the series (Pereira et al., 2021) also incorporates 9 benchmarks from Rosetta Code, a repository of even smaller and simpler code snippets such as Fibonacci, Ackermann, or Sieve of Eratosthenes (typically under 50 lines of code). Because Rosetta Code benchmarks are strictly smaller in scope and size than the CLBG benchmarks, we exclude them from consideration in this paper.
@@ 2.2.1. Programming Language versus Implementation
2.2.2. Quality of Benchmark Implementations
While benchmark implementations are claimed to employ the “exact same algorithm” across languages (Pereira et al., 2021, §1), this is in fact not the case. Benchmark implementations have highly varied levels of parallelism and CPU usage, varying degrees of use of third-party libraries, and non-uniform use of vector instructions. These important differences are not properties of the languages themselves, and their effect on performance and energy consumption must be accounted for.
@@ 2.2.3. Apparent Anomalies
Some results reported by Pereira et al. are counter-intuitive, and are presented without investigation or explanation. C++ is reported as being 34% less energy efficient and 56% slower than C. Since C++ is approximately a superset of C, and both share the same compiler, optimizations, and code generation backend, we would expect identical energy and runtime performance. Similarly, TypeScript is reported as being 4.8×4.8$\times$4.8 × less energy efficient and 7.1×7.1$\times$7.1 × slower than JavaScript. TypeScript is a strict superset of JavaScript: any valid JavaScript program is a valid TypeScript program. The same Node.js runtime is used for both languages. The compilation process for TypeScript may insert or modify code to support older JavaScript standards, but we do not expect this to result in any significant performance overhead. Java, C#, and Go are reported as 1.89×1.89$\times$1.89 ×, 3.14×3.14$\times$3.14 ×, and 2.83×2.83$\times$2.83 × slower than C, respectively. These numbers are unexpectedly high as these implementations are known to be highly optimized (Lion et al., 2022). We expect low overhead from runtime garbage collection costs, plus initial just-in-time compilation overhead for Java and C#. Lua and TypeScript both stand out as exhibiting significantly lower normalized energy consumption than normalized execution time; this is explained by the fact that nearly all of the implementations of Lua and TypeScript benchmarks are sequential, while nearly all implementations in other languages are parallelized; Section 4.4.1 characterizes the impact of the number of active cores on energy consumption.
@@ 2.2.4. Memory Metrics
Finally, total memory usage is measured by summing each measurement of RSS taken using Python’s memory_profiler, which is sampled at a frequency of ten times per second. This metric is directly proportional and thus a proxy for the benchmark’s execution time, which is itself directly proportional to energy consumption.
@@ 2.2.5. Language Implementation Specifics
Most language implementations have an initial cost to import and set up necessary in-memory data structures. Language implementations using a just-in-time compiler also require an initial warmup period (Traini et al., 2023; Barrett et al., 2017). These start-up costs are amplified by the benchmarks’ short runtimes, which in some cases run for less than a fraction of a second. For example, measuring only the first iteration of a short-lived Java benchmark may not be indicative of Java’s overall performance. Section 4.2.1 shows that this first iteration can be up to 3×3$\times$3 × slower than subsequent ones. Garbage collection also can have a significant impact on both runtime and memory usage (Hertz and Berger, 2005), and can be fine-tuned to obtain better performance.
@@ 3. Methodology
Table 2. Versions of compilers and runtimes used (§3.1). All are most recent as of September 2024.
@@ 3.1. Setup
We conduct experiments in this paper using a server equipped with two Intel Xeon Gold 6430 processors totaling 128 logical cores and 128GB of memory, running Linux version 6.8.0-45-generic. Table 2 details specific compiler and runtime versions. Experiments in both this paper and Pereira et al.’s use a server architecture.
Benchmark runs are isolated in Docker containers to ensure reproducibility. Docker introduces negligible performance overhead (Felter et al., 2015), which we confirm for the experiments detailed in this paper by obtaining equivalent results with and without Docker. Docker also provides easy access to control groups, which makes it possible to place restrictions on CPU usage or to pin programs to execute on specific cores. Section 4.3 details how we eliminate the effect of varying concurrency in different benchmark implementations by limiting benchmarks to execute on a single core.
@@ 3.2.2. Performance counters
There are multiple ways to measure average core usage while a program is running. For simplicity, we leverage performance counters provided by the operating system. The Linux kernel exposes the task-clock software event counter, which aggregates in nanoseconds the time spent on all processor cores. To obtain average core usage, we divide this counter’s value by the total runtime of the program.
3.3. Computer Language Benchmark Game
Table 3. Descriptions of benchmarks from the Computer Language Benchmark Game (§3.3), along with the range of lines of code (LoC) across languages for each benchmark as reported by cloc.
@@ 3.3. Computer Language Benchmark Game
The Computer Language Benchmark Game (CLBG) (Gouy, 2024) is a collection of small programs implemented in many different programming languages (22 to 287 lines of code). We use the same benchmark implementations as Pereira et al. (Couto and Pereira, 2017), which is in effect a snapshot of the fastest versions of the benchmarks as available on the CLBG repository at the time of the original work. Table 3 provides a brief description of each benchmark. We limit our analysis to eleven languages across thirteen implementations: C, C#, C++, Go, Java, JavaScript, Lua (Lua and LuaJIT), PHP, Python (CPython and PyPy), Rust, and TypeScript. These languages comprise 11 of the top 12 most popular general programming languages in the 2024 StackOverflow survey (Overflow, 2024) (excluding Kotlin).
We attempt to compile and run all benchmark implementations with sources present in the repository, and omit those without source code or which fail with compilation or runtime errors. After these omissions, the analysis below spans 118 out of 130 possible language implementation / benchmark pairs. Some benchmarks make use of third-party libraries: we use the default package manager’s version when available, or manually upgrade to the latest available version.
@@ 4. Causal Model for Energy Consumption in Programming Languages
This section builds piece by piece the causal model presented in its final form in Figure 2. Each subsection gradually adds nodes and edges to obtain a rich causal diagram that captures the essential factors in the relationship between the choice of programming language and energy consumption.
@@ 4.2. Programming Languages and Implementations
Figure 5 introduces the important distinction between the programming languages themselves and their implementations. Some languages have a single well-known implementation, or a blurry line between language and implementation, such as Rust or Go. Others have multiple implementations, such as C/C++ with GCC, Clang, or MSVC, Java with various JVMs, Python with CPython and PyPy, or Lua with LuaJIT.
@@
Figure 6. Programming Languages can exhibit different performance characteristics depending on the implementation used (§4.2). Here, we compare interpreters (CPython, Lua) to their JIT equivalent (PyPy, LuaJIT) on the same benchmark sources. We normalize to the interpreter results. We find around 5×5$\times$5 × and 1.25×1.25$\times$1.25 × decreased runtime and energy consumption on average for LuaJIT over Lua and PyPy over CPython, respectively. These results demonstrate the impact of choice of programming language implementation versus the language itself.
@@
Two bar charts illustrating relative runtime and energy consumption between CPython/PyPy and Lua/LuaJIT.
@@ 4.2.1. Just-in-Time (JIT) Warmup
Figure 7. Measuring just the first iteration is not indicative of Java’s (OpenJDK) performance for long-lived applications (§4.2.1). We gradually increase the number of iterations, observing decreased per-iteration runtime on several benchmarks, including (a) mandelbrot and (b) fannkuch-redux shown above. This effect could be due to startup and cleanup time and/or a JIT warmup effect. The curve flattens as we average over more iterations.
@@
Factors other than JIT warmup may also be affecting the very first iteration of those short-lived benchmarks, such as bytecode interpretation or memory and garbage collection initialization. To test this, we wrap each Java benchmark in a loop and gradually increase the number of in-process iterations. Figure 7 shows that for the mandelbrot and fannkuch-redux benchmarks, the first iteration’s runtime is 2.5×2.5$\times$2.5 × and 1.5×1.5$\times$1.5 × slower respectively than when averaging over 15 iterations. Generally, we see that for these and other benchmarks, averaging over ten or more iterations amortizes the first iteration’s impact. Applying this method over every benchmark, we observe 80% improved energy efficiency and runtime performance versus using only the first iteration.

### v2

@@ 1. Introduction
The acceleration of climate change due to use of fossil fuels has driven an increased focus on efforts to decrease both the energy consumption and carbon footprint of computer systems (Hanafy et al., 2024b, a; Radovanovic et al., 2021). In 2018, an estimated 1% of total global energy consumption was attributed to datacenters alone (Masanet et al., 2020). Modern machine learning workloads—especially training—can generate hundreds of tons of CO2 emissions (Strubell et al., 2019; Patterson et al., 2021). For example, Meta reports that training the Llama2 large language model generated an estimated 539 tons of CO2 emissions (GenAI, 2023).
@@
These studies have received wide attention, both in academic and industrial circles. They have been collectively cited nearly 700 times per Google Scholar (as of March 2025). The results—especially the ranking of programming languages—have had an unusually visible impact in industry, and are routinely quoted on social networks and in blog posts. As Figure 1 illustrates, the rankings have been cited by executives and engineers from Amazon (Miller and Lerche, 2022, 2021; Miller, 2022; Vogels, 2023), Intel (Mulhall and Collins, 2022), SAP (Muthaiah, 2022), and other companies (Suárez et al., 2021) to argue for business decisions and to advocate for a shift in programming languages with an eye towards sustainability. Often, these rankings have been harnessed to support the adoption of Rust, which ranks as one of the most energy-efficient languages while providing safety guarantees that languages like C and C++ lack.
@@
Our model specifically identifies the number of active cores and memory activity as the key contributors to power draw and consequently energy consumption. It highlights the importance of both programming language and application implementations: language implementation decisions can add execution time overheads including garbage collection or just-in-time warmup, while application implementations dictate the level of parallelism and overall performance of the program. These varying language and application implementation characteristics are the primary drivers of the anomalies we identify in prior work.
Our experimental approach incorporates several methodological improvements, correcting technical errors in prior work that led to negative or otherwise incorrect energy readings. We present an improved methodology based on hardware performance counter data that accurately quantifies average core usage and memory activity. When controlling for confounds like varying CPU utilization and normalizing by time, we conclude that energy consumed is directly proportional to execution time, and independent of the choice of programming language. These results suggest that, to minimize energy consumption, programmers should focus primarily on optimizing performance.
@@ 2.1. Overview
The line of work that is the point of departure for this paper explores the relationship between the choice of programming language and energy efficiency (Couto et al., 2017; Pereira et al., 2017, 2021). These papers rank programming languages based on energy consumption, execution time, and memory usage, and attempt to find associations between these metrics. In the remainder of this paper, we refer to these papers collectively as “Pereira et al.”.
Table 1. Partial results from Pereira et al. (§2.1), showing average execution time and energy consumption for the languages discussed in this paper, normalized to C. Boldface denotes anomalous results that Sections 2.2, 4.2, and 4.5 address.
@@
In essence, these studies compare the execution time, energy consumption and memory usage of benchmark implementations in 27 various programming languages. Table 1 presents some of the main results from Pereira et al., limited to the languages discussed in this paper, with anomalous results highlighted in boldface (Section 2.2 discusses this selection in detail). Their experiments leverage Intel’s Running Average Power Limit (RAPL) interface to measure energy consumption, and GNU time or Python’s memory_profiler to measure peak or “total” memory usage, respectively.
The programs used for comparison are from the Computer Language Benchmark Game (Gouy, 2024) (CLBG), a corpus of small benchmark implementations in various languages (22 to 287 lines of code). The most recent paper in the series (Pereira et al., 2021) also incorporates 9 benchmarks from Rosetta Code, a repository of even smaller and simpler code snippets such as Fibonacci, Ackermann, or Sieve of Eratosthenes (typically under 50 lines of code). Because Rosetta Code benchmarks are strictly smaller in scope and size than the CLBG benchmarks, we exclude them from consideration in this paper.
@@ 2.2.1. Programming Language versus Implementation
2.2.2. Selection of Benchmark Implementations
While benchmark implementations are claimed to employ the “exact same algorithm” across languages (Pereira et al., 2021, §1), this is in fact not the case. Benchmark implementations notably have highly varied levels of parallelism and CPU usage, varying degrees of use of third-party libraries, and non-uniform use of vector instructions. These important differences are not properties of the languages themselves, and their effect on performance and energy consumption must be accounted for.
@@ 2.2.3. Apparent Anomalies
Some results reported by Pereira et al. are counter-intuitive, and are presented without investigation or explanation. C++ is reported as being 34% less energy efficient and 56% slower than C. Since C++ is approximately a superset of C, and both share the same compiler, optimizations, and code generation backend, we would expect identical energy consumption and execution time. Similarly, TypeScript is reported as being 4.8×4.8$\times$4.8 × less energy efficient and 7.1×7.1$\times$7.1 × slower than JavaScript. TypeScript is a strict superset of JavaScript: any valid JavaScript program is a valid TypeScript program. The same Node.js runtime system is used for both languages. The compilation process for TypeScript may insert or modify code to support older JavaScript standards, but we do not expect this to result in any significant performance overhead. Java, C#, and Go are reported as 1.89×1.89$\times$1.89 ×, 3.14×3.14$\times$3.14 ×, and 2.83×2.83$\times$2.83 × slower than C, respectively. These numbers are unexpectedly high as these implementations are known to be highly optimized (Lion et al., 2022). We expect low overhead from garbage collection costs, plus initial just-in-time compilation overhead for Java and C#. Lua and TypeScript both stand out as exhibiting significantly lower normalized energy consumption than normalized execution time; this is explained by the fact that nearly all of the Lua and TypeScript benchmark implementations are sequential, while the implementations of these benchmarks in other languages are parallelized. Section 4.4.1 characterizes the impact of the number of active cores on energy consumption.
@@ 2.2.4. Memory Metrics
Total memory usage is measured by summing each measurement of RSS taken using Python’s memory_profiler, which is sampled at a frequency of ten times per second. This metric is directly proportional to and thus a proxy for the benchmark’s execution time, which is itself directly proportional to energy consumption.
@@ 2.2.5. Language Implementation Specifics
Most language implementations have an initial cost to import and set up necessary in-memory data structures. Language implementations using a just-in-time compiler also require an initial warmup period (Traini et al., 2023; Barrett et al., 2017). These start-up costs are amplified by the benchmarks’ short execution times, which in some cases run for less than a fraction of a second. For example, measuring only the first iteration of a short-lived Java benchmark may not be indicative of Java’s overall performance. Section 4.2.1 shows that this first iteration can be up to 3×3$\times$3 × slower than subsequent ones. Garbage collection also can have a significant impact on both execution time and memory usage (Hertz and Berger, 2005), and can be fine-tuned to obtain better performance.
@@ 3. Methodology
Table 2. Versions of compilers and runtime systems used (§3.1). All are most recent as of September 2024.
@@ 3.1. Setup
We conduct experiments in this paper using a server equipped with two Intel Xeon Gold 6430 processors totaling 128 logical cores and 128GB of memory, running Linux version 6.8.0-45-generic. Table 2 details specific compiler and runtime system versions. Experiments in both this paper and Pereira et al.’s use a server architecture.
Benchmark runs are isolated in Docker containers to ensure reproducibility. Docker introduces negligible performance overhead (Felter et al., 2015), which we confirm for the experiments detailed in this paper by obtaining equivalent results with and without Docker. Docker also provides easy access to control groups, making it possible to place restrictions on CPU usage or to pin programs to execute on specific cores. Section 4.3 details how we eliminate the effect of varying concurrency in different benchmark implementations by limiting benchmarks to execute on a single core.
@@ 3.2.2. Performance counters
There are multiple ways to measure average core usage while a program is running. For simplicity, we leverage performance counters provided by the operating system. The Linux kernel exposes the task-clock software event counter, which aggregates in nanoseconds the time spent on all processor cores. To obtain average core usage, we divide this counter’s value by the total execution time of the program.
3.3. Benchmarks
This paper primarily focuses its evaluation on the benchmarks used in the original Pereira et al. study, namely the Computer Language Benchmark Game. In some analyses, we extend that evaluation with two additional benchmark suites: SPEC CPU 2017 (Corporation, 2022) and DaCapo Chopin (Blackburn et al., 2025). The following subsections provide a brief overview of each of these benchmark suites.
3.3.1. Computer Language Benchmark Game
Table 3. Descriptions of benchmarks from the Computer Language Benchmark Game (§3.3.1), along with the range of lines of code (LoC) across languages for each benchmark as reported by cloc.
@@ 3.3.1. Computer Language Benchmark Game
The Computer Language Benchmark Game (Gouy, 2024) (CLBG) is a collection of small programs implemented in many different programming languages (22 to 287 lines of code). We use the same benchmark implementations as Pereira et al. (Couto and Pereira, 2017), which is in effect a snapshot of the fastest versions of the benchmarks as available on the CLBG repository at the time of the original work. Table 3 provides a brief description of each benchmark. We limit our analysis to eleven languages across thirteen implementations: C, C#, C++, Go, Java, JavaScript, Lua (Lua and LuaJIT), PHP, Python (CPython and PyPy), Rust, and TypeScript. These languages comprise 11 of the top 12 most popular general programming languages in the 2024 StackOverflow survey (Overflow, 2024) (excluding Kotlin).
We attempt to compile and run all benchmark implementations with sources present in the repository, and omit those without source code or which fail with compilation or run-time errors. After these omissions, the analysis below spans 118 out of 130 possible language implementation / benchmark pairs. Some benchmarks make use of third-party libraries: we use the default package manager’s version when available, or manually upgrade to the latest available version.
@@
3.3.2. SPEC CPU 2017
Table 4. SPEC CPU 2017 benchmarks (§3.3.2).
Benchmark Language(s) Description
600.perlbench_s C Perl interpreter
602.gcc_s C GNU C compiler
605.mcf_s C Route planning
620.omnetpp_s C++ Discrete-event simulation
623.xalancbmk_s C++ XML to HTML conversion
625.x264_s C Video compression
631.deepsjeng_s C++ Alpha-beta tree search
641.leela_s C++ Monte Carlo tree search
648.exchange2_s Fortran Recursive solution generator
657.xz_s C General data compression
603.bwaves_s Fortran Explosion modeling
607.cactuBSSN_s C++, C, Fortran Physics: relativity
619.lbm_s C Fluid dynamics
621.wrf_s Fortran, C Weather forecasting
627.cam4_s Fortran, C Atmosphere modeling
628.pop2_s Fortran, C Wide-scale ocean modeling
638.imagick_s C Image manipulation
644.nab_s C Molecular dynamics
649.fotonik3d_s Fortran Computational electromagnetics
654.roms_s Fortran Regional ocean modeling
Table 5. DaCapo benchmarks (§3.3.3).
Benchmark Description
avrora AVR controller simulation
batik Render SVG files
biojava Protein sequence analysis
cassandra NoSQL database
eclipse IDE workload
fop Render PDF files
graphchi ALS matrix factorization
h2 SQL database
h2o Machine learning
jme Render video game frames
jython Python implementation
kafka Stream processing
luindex Text indexing
lusearch Text search
pmd Static analysis
spring Web framework
sunflow Render photorealistic images
tomcat Web server
tradebeans Stock market simulation
tradesoap Stock market simulation
xalan XML transformation
zxing Barcode reader
SPEC CPU 2017 (Corporation, 2022) is a widely used benchmark suite for evaluating systems on C, C++ and Fortran code. The suite consists of large real-world programs and workloads, ranging between 1k and 1.3M lines of code. We use the latest available version (1.1.9) and run all benchmarks in the “speed” category using the reference input set. We omit a single benchmark, 627.cam4_s, as it did not successfully run on our platform. Table 4 provides a brief description of each benchmark.
3.3.3. DaCapo
DaCapo (Blackburn et al., 2025) is a widely used benchmark suite for Java applications. The suite consists of large real-world programs and workloads, ranging between 25k and 6M lines of code. We use the latest available version (Chopin) and run all benchmarks using the default input set. Table 5 provides a brief description of each benchmark.
@@ 4. Causal Model for Energy Consumption in Programming Languages
This section builds piece by piece the causal model presented as a diagram in its final form in Figure 2. Causal diagrams illustrate how different factors influence each other, with arrows indicating cause-to-effect relationships (Pearl, 2009). We additionally use gray arrows to represent comparatively weaker relationships, as detailed in the following sections. Each subsection below gradually adds nodes and edges to the diagram in order to obtain a rich causal model that captures the essential factors in the relationship between the choice of programming language and energy consumption. Figures 4, 5, 9, 11, and 13 build upon each other, with new elements in orange.
@@ 4.2. Programming Languages and Implementations
Figure 6. Execution time for the ten longest-running pyperformance benchmarks in the five last Python releases, normalized to CPython 3.9 (§4.2). The thicker gray line indicates $y=1$. Most benchmarks show improvements over time, with a $40\%$ speedup on average for the entire suite between CPython 3.9 and 3.13.
[…]
A line graph with 10 lines, one for each benchmark. 9 out of 10 benchmarks end under the $y=1$ line for CPython 3.13.
Figure 5 introduces the important distinction between the programming languages themselves and their implementations. Some languages have a single well-known implementation, or a blurry line between language and implementation, such as Rust or Go. Others have multiple implementations, such as C/C++ with GCC, Clang, or MSVC, Java with various JVMs, Python with CPython and PyPy, or Lua with LuaJIT. Finally, there can be significant differences between specific versions of the same implementation. Our model treats those different versions as different implementations altogether. For example, we test the pyperformance benchmark suite (pyperformance, 2025) across CPython versions 3.9 through 3.13. Figure 6 shows overall improvements for the ten longest-running benchmarks. Over the entire suite, we observe a $40\%$ speedup using CPython 3.13 over 3.9.
@@
Figure 7. Programming Languages can exhibit different performance characteristics depending on the implementation used (§4.2). Here, we compare interpreters (CPython, Lua) to their JIT equivalent (PyPy, LuaJIT) on the same benchmark sources. We normalize to the interpreter results. We find around 5×5$\times$5 × and 1.25×1.25$\times$1.25 × decreased execution time and energy consumption on average for LuaJIT over Lua and PyPy over CPython, respectively. These results demonstrate the impact of choice of programming language implementation versus the language itself.
@@
Two bar charts illustrating relative execution time and energy consumption between CPython/PyPy and Lua/LuaJIT.
@@ 4.2.1. Just-in-Time (JIT) Warmup
Figure 8. Measuring just the first iteration is not indicative of Java’s (OpenJDK) performance for long-lived applications (§4.2.1). We gradually increase the number of iterations, observing decreased per-iteration execution time on several benchmarks, including (a) mandelbrot and (b) fannkuch-redux shown above. This effect could be due to startup and cleanup time and/or a JIT warmup effect. The curve flattens as we average over more iterations.
@@
Factors other than JIT warmup may also be affecting the very first iteration of those short-lived benchmarks, such as bytecode interpretation or memory and garbage collection initialization. To test this, we wrap each Java benchmark in a loop and gradually increase the number of in-process iterations. Figure 8 shows that for the mandelbrot and fannkuch-redux benchmarks, the first iteration’s execution time is 2.5×2.5$\times$2.5 × and 1.5×1.5$\times$1.5 × slower respectively than when averaging over 15 iterations. Generally, we see that for these and other benchmarks, averaging over ten or more iterations amortizes the first iteration’s impact. Applying this method over every benchmark, we observe 80% improved energy efficiency and execution time versus using only the first iteration.